Suppose the real rate is 2.25 percent and the inflation rate is 3.2 percent. What rate would you expect to see on a Treasury bill?
Answer #1

(1+Real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.0225)=(1+nominal rate)/(1+0.032)

1+nominal rate=(1.0225*1.032)

nominal rate=(1.0225*1.032)-1

=5.522%